A triangle has a height of 6cm and a base of 8 cm. If the height and the base are both decreased by the same amount, the area of the new triangle is 20 cm squared. what is the base and height of the new triangle?
Answer by mananth(16946)

height 6 cm
base 8 cm
decreased by x cm
6-x & 8-x will be the new height and base
Arae of triangle = 1/2 *(6-x)(8-x)=20
(6-x)(8-x)=20*2
48-14x+x^2=40
x^2-14x +8=0
use quadratic formula
discriminant = 164
x1= (14+sqrt164)/2
x1=(14+12.8)/2
=26.8/2
23.4 ( not possible)
x2= 14-sqrt164)/2
x2=14-12.8)/2
x2=1.2/2
x2=0.6 cm
...
The new dimensions are height = 5.4, base = 7.4
...
m.ananth@hotmail.ca